OpenStudy (anonymous):

A triangle that has a hypotenuse of length 14.14 cm has legs that are ______ cm long. Round your answer to the nearest whole centimeter. (Enter the number only.)

OpenStudy (anonymous):

Is there a picture that goes along with this problem?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

Because if only the hypotenuse is given, the legs of the right triangle can be many different lengths.

OpenStudy (anonymous):

Is there any other info given?

OpenStudy (anonymous):

Oh, according to what I read in the question, this is an isosceles right triangle they're talking about.

OpenStudy (anonymous):

So:

OpenStudy (anonymous):

\[\large 14.14^2 = x^2 + x^2 = 2x^2\] \[\large \frac {14.14^2}{2} = x^2\] \[\large \sqrt {\frac{14.14^2}{2}} = x\]

OpenStudy (anonymous):

Try that...

OpenStudy (anonymous):

I walked away sorry